# Privacy-Preserving In-Context Learning with Differentially Private Few-Shot Generation

### v1

@@ 5.3 Ablation Studies
In this section, we conduct ablation studies on various number of shots, LM sizes, and DP mechanisms in Algorithm 1. Due to the sheer amount of experiments with 5 rounds each, we fix the dataset to DBPedia in our ablation studies.
5.3.1 Varying number of shots
In Table 2, we present the results with various number of shots for ICL. $\epsilon=0$ represents the performance of synthetic few-shot demonstrations that are generated with purely the instruction without any private data (i.e. running only line 1 auto-regressively in Algorithm 1). $\epsilon=4$ presents the performance of our private solution in Algorithm 1. We also separately present two non-private performances; (i) running Algorithm 1 with $\sigma=0$ and (ii) n-shot demonstrations randomly selected from the private dataset.
We observe that our approach consistently improves the fully private $\epsilon=0$ performance for all the n-shot cases. Furthermore, our approach is competitive with both non-private performances for 1, 2, 4-shot cases. An interesting observation is that for 8-shot and 12-shot cases, randomly picking the samples from the private dataset provides better performance. Reducing the noise does not bridge this gap because $\epsilon=4$ and $\epsilon=\infty$ ($\sigma=0$ in Alg. 1) have competitive performances. Therefore, this shows that Algorithm 1 is open to improvements for larger n-shot scenarios.
Table 2: ICL performance on the test set of DBPedia dataset with various number of shots. Mean and standard deviation of accuracy is reported over 5 runs with different random seeds. $\epsilon=0$ is the fully private solution that uses purely the public instruction to generate synthetic few-shot demonstrations for ICL. $\epsilon=4$ is based on the generations of our Algorithm 1 with private data for ICL. $\epsilon=\infty$ ($\sigma=0$ in Alg. 1) uses the generations of our Algorithm 1 with $\sigma=0$ and $\epsilon=\infty$ (random) uses randomly chosen n-shot samples from the private data. Our approach substantially improves the performance of fully private solution and performs competitively with non-private solutions for 1, 2, and 4-shots. For 8 and 12-shots, demonstrations directly with private samples perform better, suggesting potential improvements of our approach for large n-shot scenarios.
@@ 5.3.1 Varying number of shots
5.3.2 Varying the model size
We next consider various model sizes (Ada, Babbage, Curie, and Davinci GPT-3 models) and present the results for ICL in Table 3. We note that generating synthetic few-shot demonstrations with purely the instruction without any private data ($\epsilon=0$ (4-shot) in Table 3) improves ICL performance as the model becomes more capable with larger sizes as expected. Our approach provides further gains on this fully private solution until the Davinci model, for which the fully private solution matches the non-private solution. This emphasizes that very large language models can at times possess the information to solve a task and while the naive 0-shot approach does not provide good performance, one can extract useful information from the model’s existing knowledge from the pre-training and use it in ICL to substantially improve the performance.
Table 3: ICL performance on the test set of DBPedia dataset with various model sizes. Mean and standard deviation of accuracy is reported over 5 runs with different random seeds. The column descriptions follow similarly as Table 1. Our approach substantially improves the performance of fully private solutions for Ada, Babbage, and Curie models. The most powerful Davinci model, although performs poorly in 0-shot case, can generate useful demonstrations from its existing knowledge for itself to use it in ICL and even match the non-private solution.
@@ 5.3.2 Varying the model size
5.3.3 Gaussian mechanism vs Exponential mechanism
We finally experiment with the Gaussian and Exponential mechanisms for the DP mechanism in Algorithm 1 and present the results in Table 4. Table 4 shows that instantiating Algorithm 1 with different DP mechanisms at $\epsilon\in\{1,2,4,8\}$ provides competitive performances. Here, we point out advantages and disadvantages of Gaussian and Exponential mechanisms. As stated in Appendix A, the Exponential mechanism is based on Report-Noisy-Max framework, which can only generate a token with the highest noisy probability to satisfy DP. In contrast, the Gaussian mechanism benefits from the post-processing property of DP and is suitable for more advanced sampling techniques beyond $\arg\max$ such as top-k sampling [18] or nucleus sampling [22]. On the other hand, the Exponential mechanism gives a pure DP guarantee of $\delta=0$ which can be of interest in some scenarios. However, if we insist on $\delta=0$ during the composition of the privacy loss across $T_{\max}$ iterations, our subsampled DP algorithm leads to a privacy loss that grows linearly in $T_{\max}$.
Table 4: ICL performance on the test set of DBPedia dataset comparing Gaussian and Exponential mechanisms for the DP mechanism in Algorithm 1. Mean and standard deviation of accuracy is reported over 5 runs with different random seeds. We observe in general competitive performances between the two DP mechanisms.
@@ 6 Related Work
In this work, we focus on in-context learning (ICL) framework, where few demonstrations from a dataset suffices LLMs to perform well on a downstream task without requiring any fine-tuning as recently observed in [7]. As we focus on the privacy-preserving aspect of the ICL framework, we refer the readers to the survey [12] and the references therein for an extensive study about this popular direction.
@@
In the context of privacy-preserving ICL, very recently, Panda et al. [30] propose DP inference by privately constructing consensus over an ensemble of queries with disjoint sampled demonstrations. Although this approach adheres to DP guarantees, it expends the privacy budget for each query. Hence, for a given privacy budget, their algorithm can only answer few limited number of queries where as our approach can answer unbounded number of queries.
@@
In another recent work, Duan et al. [14] propose an approach that also relies on private ensembling. However, their approach assumes access to some unlabeled public data, which is privately labeled by an ensemble of teachers with ICL via demonstrations from the private dataset. Privately labeled public data is finally used as the prompt for ICL. This is in contrast to our approach that does not require any public data. Our framework is more suitable for the scenarios where the availability of unlabeled public data that has similar distribution as the private data may be a strong assumption; examples include healthcare datasets or industrial applications.
@@
We further note that Panda et al. [30], Duan et al. [14] mainly focus on text classification tasks while our method readily extends beyond text classification as demonstrated by our experiments.
@@
Our work focuses on privacy-preserving few-shot generation for ICL without the heavy burden of private fine-tuning [42, 23]. We leverage the capabilities of LLMs [7] that can generate text that is similar to the given demonstrations and convert this process into a DP one via privately aggregating the generation probabilities obtained from disjoint subsets of demonstrations from the private data. This approach is in spirit similar to Private Aggregation of Teacher Ensembles (PATE) framework [31, 32], which produces a private model via fine-tuning with public data, which is privately labeled using an ensemble of teachers that are trained on disjoint subsets of the private data. In this line of work, SeqPATE [34] and Submix [20] extend the PATE framework to text generation via several design adaptions for text domain. On the other hand, Majmudar et al. [25] introduces DP at the decoding stage of a pre-trained LLM by combining the prediction vector of the pre-trained LLM and uniform distribution for DP decoding. Note that these private inference methods still require (non-private) training the model with the private data to adapt to the specific data domain.
@@
Our work can be seen as an instantiation of the more general synthetic text generation with privacy framework. As opposed to the approaches that require private fine-tuning [43, 26, 29, 8] in this direction, we introduce a light-weight approach that can utilize LLMs without needing the computationally prohibitive fine-tuning. We point out that our work is efficient and effective in ICL framework where few-shot demonstrations suffice. However, we believe that private fine-tuning would be more effective within applications that require generating large volume of synthetic data.
Besides, another line of work [19, 41, 13] focuses on sanitizing user texts locally before releasing them to the server based on metric local differential privacy [9]. Such methods usually incur huge overheads to the utility of the sanitized text. The utility can be improved by incorporating inherent properties of language such as semantic similarity [10], context [3] and syntax [2, 29].
@@ 7 Conclusion and Future Work
The main weakness of our Algorithm 1 is repeating line 1 and 1, i.e. resampling the demonstrations from the private dataset for each token generation. It would be more natural to fix the demonstrations from the private dataset and generate one synthetic sample all at once. Unfortunately this does not benefit from the amplification of privacy by subsampling theorem and requires much higher levels of noise additions in line 13. We believe that our algorithm can be improved on this front in future work. Furthermore, as stated in Section 5.3.3, Gaussian mechanism benefits from the post-processing of DP and more advanced sampling techniques for token generation can be used instead of the basic $\arg\max$ sampling. We leave the exploration of this to future work.
@@
The following theorem concerning the privacy curve of Gaussian mechanism is due to Balle and Wang [2018].
@@
The advantage of working with privacy curves of DP mechanisms is that it can be used to get tighter guarantees on the composition compared to the advanced composition theorems [Dwork and Roth, 2014]. In particular, one can do numerical composition of privacy curves, which gives the tightest guarantees.
Theorem A.3 (Gopi et al. [2021]).
@@
The work of Gopi et al. [2021] gives privacy curves for composition of several standard mechanisms such as Gaussian mechanism, Laplace mechanism, and subsampled Gaussian mechanism.
The variant of our algorithm using exponential noise is based on the Report Noisy Max with Exponential noise mechanism [Dwork and Roth, 2014]. The exponential distribution, denoted by $\mathrm{Expo}(\lambda)$, has the density function $f(x;\lambda)=\lambda\exp(-\lambda x)\mathbf{1}_{x\geq 0}$. The mean of $\mathrm{Expo}(\lambda)$ is $1/\lambda$. This algorithm adds to every possible output noise drawn from $\mathrm{Expo}(\frac{\epsilon}{2\Delta})$ and reports the resulting arg-max, where $\Delta$ is the $\ell_{\infty}$-sensitivity of the resulting output (or more generally scoring function used to select the outputs). It was shown in Ding et al. [2021] that the Report Noisy Max with Exponential noise is equivalent to the Permute-and-Flip mechanism, which was proposed by McKenna and Sheldon [2020] as an improvement over the popular exponential mechanism.
@@
Unlike the Gaussian mechanism, this mechanism gives pure DP guarantee; i.e., $\delta=0$. Finally, we need the following amplification of privacy by subsampling theorem due to Ullman [2017], Balle et al. [2018].
@@
Consider the Gaussian mechanism first. Fix one iteration of the algorithm and let us bound the privacy loss. In line 1, $\ell_{2}$-sensitivity of the term $\left(\sum_{i=1}^{M}p_{\text{priv}}^{i}\right)$ is at most $\sqrt{2}$. As we are adding noise sampled from $\mathcal{N}(0,2\sigma^{2}\mathbf{I})$, it follows that our algorithm has the same privacy curve as given in Theorem A.2. Further, note that in each iteration we draw $MN$ samples from a dataset of size $|\mathcal{D}_{\text{priv}}|$. Thus the effective privacy curve of our algorithm per iteration is given by the privacy curve of subsampled Gaussian mechanism, which is calculated in Gopi et al. [2021]. Thus, the output of our algorithm $\widetilde{p}_{\text{priv}}$ on line 1 satisfies $(\epsilon,\delta)$-DP guarantees, and further operations done by our algorithm on the $\widetilde{p}_{\text{priv}}$ does not violate the privacy guarantees due to the post-processing property of DP [Dwork and Roth, 2014]. To get an upper bound on the privacy loss across all $T_{max}$ iterations, we appeal to Theorem A.3 for composing of $T_{\max}$ privacy curves of subsampled Gaussian mechanisms [Gopi et al., 2021].
@@
As Theorem A.3 is based on numerical composition, the final privacy parameters of our algorithms do not have closed forms. Readers can refer to Abadi et al. [2016] to get an analytical expression of the privacy loss, although it does not provide the tight upper bound.
@@ Appendix B Prompt construction functions for each task in Section 5
In Table 5 and 6 we provide our prompt construction function $PB(\text{instruction},\mathcal{D}^{y},y)$ for each task in Section 5.
@@ Appendix C Datasets used in Section 5
The AG News (AG’s News Corpus) dataset [Zhang et al., 2015] consists of news articles belonging one of the 4 labels (World, Sports, Business, Sci/Tech). The AG News has 30,000 training and 1,900 test samples per class.
@@
The Text REtrieval Conference (TREC) question classification dataset [Voorhees and Tice, 2000] consists of questions that have answers belonging one of the 6 answer types (Number, Location, Person, Description, Entity, Abbreviation). The TREC has in total 5,500 training and 500 test samples that are non-uniform across the labels.
@@
The DBpedia ontology classification dataset [Zhang et al., 2015] consists of contents belonging one of the 14 topics (Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, Book). The DBpedia has 40,000 training samples and 5,000 testing sample per class.
@@
MIT Movies trivia10k13 dataset [Liu et al., 2012] consists of movie reviews that have slots (movie genre (MIT-G) and director name (MIT-D)) to be used for information extraction task. MIT-G has 2,953 training samples and 100 test samples whereas MIT-D has 1,561 training samples and 100 test samples.
Appendix D Prompt format during ICL following prior work Zhao et al. [2021]
While we use the same prompt format during ICL following prior work by Zhao et al. [2021], we present them here in Table 7 and 8 for the convenience of the reader.
Table 7: The prompts used during ICL for text classification tasks, taken from Table 5 of Zhao et al. [2021].
@@
Table 8: The prompts used during ICL for information extraction tasks, taken from Table 6 of Zhao et al. [2021].

### v2

@@ 5.3 Ablation Studies
In this section, we conduct ablation studies on various number of shots, LM sizes, and DP mechanisms in Alg. 1. Due to the sheer amount of experiments with 5 rounds each, we fix the dataset to DBPedia.
Varying number of shots. In Tab. 2, we present the results with various number of shots for ICL. $\epsilon=0$ represents the performance of synthetic few-shot demonstrations that are generated with purely the instruction without any private data (i.e. running only line 1 in Alg. 1). $\epsilon=4$ presents the performance of our private solution in Alg. 1. We separately present two non-private performances; (i) running Alg. 1 with $\sigma=0$ and (ii) n-shot samples randomly selected from the private dataset.
Table 2: ICL performance on the test set of DBPedia dataset with various number of shots. $\epsilon=0$ is the fully private solution that uses purely the public instruction to generate synthetic few-shot demonstrations for ICL. $\epsilon=4$ is based on the generations of our Alg. 1 with private data for ICL. $\epsilon=\infty$ ($\sigma=0$ in Alg. 1) uses the generations of our Alg. 1 with $\sigma=0$ and $\epsilon=\infty$ (random) uses randomly chosen n-shot samples from the private data. Our approach substantially improves the performance of fully private solution and performs competitive with non-private solutions for 1, 2, and 4-shots. For 8 and 12-shots, demonstrations directly with private samples perform better, suggesting potential improvements of our approach for large n-shot scenarios.
@@
We observe that our approach consistently improves the fully private $\epsilon=0$ performance for all the n-shot cases. Furthermore, our approach is competitive with both non-private performances for 1, 2, 4-shot cases. An interesting observation is that for 8-shot and 12-shot cases, randomly picking the samples from the private dataset yields better performance. Reducing the noise does not bridge this gap because $\epsilon=4$ and $\epsilon=\infty$ ($\sigma=0$ in Alg. 1) have competitive performances. Therefore, this shows that Alg. 1 is open to improvements for larger n-shot scenarios.
Varying the model size. We next consider various model sizes (Ada, Babbage, Curie, and Davinci GPT-3 models) and present the results for ICL in Tab. 3. We note that generating synthetic few-shot demonstrations with purely the instruction without any private data ($\epsilon=0$ (4-shot) in Tab. 3) improves ICL performance as the model becomes more capable with larger sizes as expected. Our approach provides further gains on this fully private solution until the Davinci model, for which the fully private solution matches the non-private solution. This emphasizes that very large language models can at times possess the information to solve a task and while the naive 0-shot approach does not provide good performance, one can extract useful information from the model’s existing knowledge from the pre-training and use it in ICL to substantially improve the performance.
Table 3: ICL performance on the test set of DBPedia dataset with various model sizes. The column descriptions follow similarly as Tab. 1. Our approach substantially improves the performance of fully private solutions for Ada, Babbage, and Curie models. The most powerful Davinci model, although performs poorly in 0-shot case, can generate useful demonstrations from its existing knowledge for itself to use it in ICL and even match the non-private solution.
@@
Gaussian mechanism vs Exponential mechanism. We finally experiment with the Gaussian and Exponential mechanisms for the DP mechanism in Alg. 1 and present the results in Tab. 4. Tab. 4 shows that instantiating Alg. 1 with different DP mechanisms at $\epsilon\in\{1,2,4,8\}$ provides competitive performances. Here, we point out advantages and disadvantages of Gaussian and Exponential mechanisms. As stated in Appendix A, the Exponential mechanism is based on Report-Noisy-Max framework, which can only generate a token with the highest noisy probability to satisfy DP. In contrast, the Gaussian mechanism benefits from the post-processing property of DP and is suitable for more advanced sampling techniques beyond $\arg\max$ such as nucleus sampling (Holtzman et al., 2020). Finally, Exponential mechanism gives a pure DP guarantee of $\delta=0$ which can be of interest in some scenarios. However, if we insist on $\delta=0$ during the composition of the privacy loss across $T_{\max}$ iterations, our subsampled DP algorithm leads to a privacy loss that grows linearly in $T_{\max}$.
Table 4: ICL performance on the test set of DBPedia dataset comparing Gaussian and Exponential mechanisms for the DP mechanism in Alg. 1. We observe in general competitive performances between the two DP mechanisms.
@@
Furthermore, we perform an empirical privacy analysis in Appendix F via membership inference attacks (MIA) (Shokri et al., 2017; Duan et al., 2023). The result shows that while the non-private ICL incurs a significant membership privacy leakage, the DP few-shot samples generated by Alg. 1 effectively reduce MIA to random guess. Finally, we discuss the monetary cost of Alg. 1 in Appendix G and provide randomly chosen generated synthetic samples in Appendix H.
@@ 6 Related Work
In this work, we focus on in-context learning (ICL) framework, where few demonstrations from a dataset suffices LLMs to perform well on a downstream task without requiring any fine-tuning as recently observed in Brown et al. (2020). As we focus on the privacy-preserving aspect of the ICL framework, we refer the readers to the survey (Dong et al., 2023) and the references therein for an extensive study about this popular direction.
@@
In the context of privacy-preserving ICL, very recently, Wu et al. (2024) propose DP inference by privately constructing consensus over an ensemble of queries with disjoint sampled demonstrations. Although this approach adheres to DP guarantees, it expends the privacy budget for each query. Hence, for a given privacy budget, their algorithm can only answer few limited number of queries where as our approach can answer unbounded number of queries.
@@
In another recent work, Duan et al. (2023) propose an approach that also relies on private ensembling. However, their approach assumes access to some unlabeled public data, which is privately labeled by an ensemble of teachers with ICL via demonstrations from the private dataset. Privately labeled public data is finally used as the prompt for ICL. This is in contrast to our approach that does not require any public data. Our framework is more suitable for the scenarios where the availability of unlabeled public data that has similar distribution as the private data may be a strong assumption; examples include health care datasets or industrial applications.
@@
We further note that Wu et al. (2024), Duan et al. (2023) mainly focus on text classification tasks while our method readily extends beyond text classification as demonstrated by our experiments.
@@
Our work focuses on privacy-preserving few-shot generation for ICL without the heavy burden of private fine-tuning (Yu et al., 2022; Li et al., 2022). We leverage the capabilities of LLMs (Brown et al., 2020) that can generate text that is similar to the given demonstrations and convert this process into a DP one via privately aggregating the generation probabilities obtained from disjoint subsets of demonstrations from the private data. This approach is in spirit similar to Private Aggregation of Teacher Ensembles (PATE) framework (Papernot et al., 2017; 2018), which produces a private model via fine-tuning with public data, which is privately labeled using an ensemble of teachers that are trained on disjoint subsets of the private data. In this line of work, SeqPATE (Tian et al., 2022) and Submix (Ginart et al., 2022) extend the PATE framework to text generation via several design adaptions for text domain. On the other hand, Majmudar et al. (2022) introduces DP at the decoding stage of a pre-trained LLM by combining the prediction vector of the pre-trained LLM and uniform distribution for DP decoding. Note that these private inference methods still require (non-private) training the model with the private data to adapt to the specific data domain.
@@
Our work can be seen as an instantiation of the more general synthetic text generation with privacy framework. As opposed to the approaches that require private fine-tuning (Yue et al., 2023; Mattern et al., 2022a; Mireshghallah et al., 2023; Carranza et al., 2023) in this direction, we introduce a light-weight approach that can utilize LLMs without needing the computationally prohibitive fine-tuning. We point out that our work is efficient and effective in ICL framework where few-shot demonstrations suffice. However, we believe that private fine-tuning would be more effective within applications that require generating large volume of synthetic data.
We finally review an active area of research on the generation of sanitized text documents. In this domain, Feyisetan et al. (2020); Xu et al. (2020); Carvalho et al. (2021); Du et al. (2023) add noise at the word level and use metric local differential privacy (Chatzikokolakis et al., 2013). Mattern et al. (2022b); Utpala et al. (2023) directly operate on documents using local differential privacy (LDP) (Kasiviswanathan et al., 2010; Duchi et al., 2013) where the former is based on fine-tuning for paraphrasing and the latter uses zero-shot prompting for paraphrasing followed by sanitization.
@@ 7 Conclusion and Future Work
The main weakness of our Alg. 1 is repeating line 1 and 1, i.e. resampling the demonstrations from the private dataset for each token generation. It would be more natural to fix the demonstrations from the private dataset and generate one synthetic sample all at once. Unfortunately this does not benefit from the amplification of privacy by subsampling theorem and requires much higher levels of noise additions. We believe that our algorithm can be improved on this front in future work. Furthermore, as stated in Section 5.3, Gaussian mechanism benefits from the post-processing of DP and more advanced sampling techniques for token generation can be used instead of the basic $\arg\max$ sampling. We leave the exploration of this to future work.
@@
The following theorem concerning the privacy curve of Gaussian mechanism is due to Balle & Wang (2018).
@@
The advantage of working with privacy curves of DP mechanisms is that it can be used to get tighter guarantees on the composition compared to the advanced composition theorems (Dwork & Roth, 2014). In particular, one can do numerical composition of privacy curves, which gives the tightest guarantees.
Theorem A.3 (Gopi et al. (2021)).
@@
The work of Gopi et al. (2021) gives privacy curves for composition of several standard mechanisms such as Gaussian mechanism, Laplace mechanism, and subsampled Gaussian mechanism.
The variant of our algorithm using exponential noise is based on the Report Noisy Max with Exponential noise mechanism (Dwork & Roth, 2014). The exponential distribution, denoted by $\mathrm{Expo}(\lambda)$, has the density function $f(x;\lambda)=\lambda\exp(-\lambda x)\mathbf{1}_{x\geq 0}$. The mean of $\mathrm{Expo}(\lambda)$ is $1/\lambda$. This algorithm adds to every possible output noise drawn from $\mathrm{Expo}(\frac{\epsilon}{2\Delta})$ and reports the resulting arg-max, where $\Delta$ is the $\ell_{\infty}$-sensitivity of the resulting output (or more generally scoring function used to select the outputs). It was shown in Ding et al. (2021) that the Report Noisy Max with Exponential noise is equivalent to the Permute-and-Flip mechanism, which was proposed by McKenna & Sheldon (2020) as an improvement over the popular exponential mechanism.
@@
Unlike the Gaussian mechanism, this mechanism gives pure DP guarantee; i.e., $\delta=0$. Finally, we need the following amplification of privacy by subsampling theorem due to Ullman (2017), Balle et al. (2018).
@@
Consider the Gaussian mechanism first. Fix one iteration of the algorithm and let us bound the privacy loss. In line 1, $\ell_{2}$-sensitivity of the term $\left(\sum_{i=1}^{M}p_{\text{priv}}^{i}\right)$ is at most $\sqrt{2}$. As we are adding noise sampled from $\mathcal{N}(0,2\sigma^{2}\mathbf{I})$, it follows that our algorithm has the same privacy curve as given in Theorem A.2. Further, note that in each iteration we draw $MN$ samples from a dataset of size $|\mathcal{D}_{\text{priv}}|$. Thus the effective privacy curve of our algorithm per iteration is given by the privacy curve of subsampled Gaussian mechanism, which is calculated in Gopi et al. (2021). Thus, the output of our algorithm $\widetilde{p}_{\text{priv}}$ on line 1 satisfies $(\epsilon,\delta)$-DP guarantees, and further operations done by our algorithm on the $\widetilde{p}_{\text{priv}}$ does not violate the privacy guarantees due to the post-processing property of DP (Dwork & Roth, 2014). To get an upper bound on the privacy loss across all $T_{max}$ iterations, we appeal to Theorem A.3 for composing of $T_{\max}$ privacy curves of subsampled Gaussian mechanisms (Gopi et al., 2021).
@@
As Theorem A.3 is based on numerical composition, the final privacy parameters of our algorithms do not have closed forms. Readers can refer to Abadi et al. (2016) to get an analytical expression of the privacy loss, although it does not provide the tight upper bound.
@@ Appendix B Prompt construction functions for each task in Section 5
In Tab. 5 and 6 we provide our prompt construction function $PB(\text{instruction},\mathcal{D}^{y},y)$ for each task in Section 5.
@@ Appendix C Datasets used in Section 5
The AG News (AG’s News Corpus) dataset (Zhang et al., 2015) consists of news articles belonging one of the 4 labels (World, Sports, Business, Sci/Tech). The AG News has 30,000 training and 1,900 test samples per class.
@@
The Text REtrieval Conference (TREC) question classification dataset (Voorhees & Tice, 2000) consists of questions that have answers belonging one of the 6 answer types (Number, Location, Person, Description, Entity, Abbreviation). The TREC has in total 5,500 training and 500 test samples that are non-uniform across the labels.
@@
The DBpedia ontology classification dataset (Zhang et al., 2015) consists of contents belonging one of the 14 topics (Company, School, Artist, Athlete, Politician, Transportation, Building, Nature, Village, Animal, Plant, Album, Film, Book). The DBpedia has 40,000 training samples and 5,000 testing sample per class.
@@
MIT Movies trivia10k13 dataset (Liu et al., 2012) consists of movie reviews that have slots (movie genre (MIT-G) and director name (MIT-D)) to be used for information extraction task. MIT-G has 2,953 training samples and 100 test samples whereas MIT-D has 1,561 training samples and 100 test samples.
Appendix D Prompt format during ICL following prior work Zhao et al. (2021)
While we use the same prompt format during ICL following prior work by Zhao et al. (2021), we present them here in Tab. 7 and 8 for the convenience of the reader.
Table 7: The prompts used during ICL for text classification tasks, taken from Tab. 5 of Zhao et al. (2021).
@@
Table 8: The prompts used during ICL in information extraction tasks, taken from Tab. 6 of Zhao et al. (2021).